OpenStudy (anonymous):

How do i evaluate the integral of ln(x^2+10x+25)?

OpenStudy (anonymous):

\[x^2+10x+25=(x+5)^2\] so \[\ln(x^2+10x+25)=\ln(x+5)^2=2\ln(x+5)\] and you can integrate by parts.